OpenStudy (anonymous):

NEED HELP To which family does the function y=2^x+5 belong quadratic square root exponential logarithmic

OpenStudy (anonymous):

Same question for the equation y=(x+2)^(1/2)+3 The answers can be quadratic square root exponential reciprocal I know it's not exponential

OpenStudy (jdoe0001):

\(\Large y=2^{\color{blue}{ x}}+5\impliedby hint\)

OpenStudy (anonymous):

Well, I already know what the equation is... I just don't understand how to solve :/

OpenStudy (jdoe0001):

we.. you asked to which it belongs to not to solve it

OpenStudy (jdoe0001):

seems solved already anyway

OpenStudy (anonymous):

I guess I don't understand completely what the differences are.

OpenStudy (jdoe0001):

\(\Large y=(x+2)^{\frac{1}{2}+3}?\)

OpenStudy (jdoe0001):

well.. can't be... so I assume is hmmm \(\Large y=(x+2)^{\frac{1}{2}}+3?\)

OpenStudy (anonymous):

The +3 is not connected to the exponent

OpenStudy (anonymous):

Yes

OpenStudy (jdoe0001):

\(\bf y=(x+2)^{\frac{1}{2}}+3\implies y-3=(x+2)^{\frac{1}{2}} \\ \quad \\ y-3=\sqrt{x+2}\implies (y-3)^2=(\sqrt{x+2})^2\implies (y-3)^2=x+2 \\ \quad \\ (y-3)^2-2=x\implies y^2-6y+9-2=x \\ \quad \\ y^2-6y+7=x\impliedby quadratic\)

OpenStudy (anonymous):

Thank you so much! I will write this down in my notes!! :) Is the first one B?

OpenStudy (jdoe0001):

and yes, both have an exponent, and one could say they're both exponenttials BUT the difference being, one exponent is an rational or fraction the other is not so the rational can always be expressed as a root and from there you end up with some exponent on the other side

OpenStudy (anonymous):

Oooooooh

OpenStudy (jdoe0001):

\(\Large { a^{\frac{{\color{blue} n}}{{\color{red} m}}} \implies \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}\implies a^{\frac{{\color{blue} n}}{{\color{red} m}}} \\\quad \\ % rational negative exponent a^{-\frac{{\color{blue} n}}{{\color{red} m}}} = \cfrac{1}{a^{\frac{{\color{blue} n}}{{\color{red} m}}}} \implies \cfrac{1}{\sqrt[{\color{red} m}]{a^{\color{blue} n}}}\qquad\qquad % radical denominator \cfrac{1}{\sqrt[{\color{red} m}]{a^{\color{blue} n}}}= \cfrac{1}{a^{\frac{{\color{blue} n}}{{\color{red} m}}}}\implies a^{-\frac{{\color{blue} n}}{{\color{red} m}}} \\\quad \\ }\) thus